# A Sober Look at Progress in Language Model Reasoning: Pitfalls and Paths to Reproducibility

### v1

@@ Abstract
Reasoning has emerged as the next major frontier for language models (LMs), with rapid advances from both academic and industrial labs. However, this progress often outpaces methodological rigor, with many evaluations relying on benchmarking practices that lack transparency, robustness, or statistical grounding. In this work, we conduct a comprehensive empirical study and find that current mathematical reasoning benchmarks are highly sensitive to subtle implementation choices—including decoding parameters, random seeds, prompt formatting, and even hardware and software-framework configurations. Performance gains reported in recent studies frequently hinge on unclear comparisons or unreported sources of variance. To address these issues, we propose a standardized evaluation framework with clearly defined best practices and reporting standards. Using this framework, we reassess recent methods and find that reinforcement learning (RL) approaches yield only modest improvements—far below prior claims—and are prone to overfitting, especially on small-scale benchmarks like AIME’24. In contrast, supervised finetuning (SFT) methods show consistently stronger generalization. To foster reproducibility, we release all code, prompts, and model outputs, for reasoning benchmarks, establishing more rigorous foundations for future work.
[…]
@@ 1 Introduction
Reasoning has become central to recent advances in large language models (LLMs), playing a key role in nearly all frontier systems (Jaech et al., 2024; Anthropic, 2025; OpenAI, 2025; xAI, 2025; Meta-AI, 2025; DeepMind, 2025). Recent months have seen a surge of research focused on understanding and improving LLM reasoning, accompanied by several open-source tools and training strategies (see Li et al. (2025b) for a survey). This momentum has sparked optimism that building capable, competitive reasoning models may soon be within reach.
@@
However, as evaluation practices shape the direction and perceived progress of the field (Liao et al., 2021), concerns around methodological rigor are growing. Non-reproducible or inconclusive evaluations can distort scientific understanding, misguide adoption, and skew future research priorities (Henderson et al., 2018; Marie et al., 2021; Musgrave et al., 2020; Prabhu et al., 2020; Andrychowicz et al., 2020; Colas et al., 2018). In the fast-moving area of LLM reasoning, where rapid publication cycles and benchmarking races are common, methodological shortcuts can quietly undermine progress. While concerns about reproducibility in LLM evaluations are well-documented (Reuel et al., 2024; Biderman et al., 2024), their persistence—especially in reasoning—calls for renewed scrutiny and higher standards.
@@
Motivated by a growing number of inconsistent empirical claims across the reasoning landscape, we conduct a rigorous investigation into the current state of reasoning benchmarks, focusing specifically on mathematical reasoning—one of the most widely used testbeds for evaluating algorithmic advances in this space (HuggingFaceH4, 2024; AI-MO, ).
@@
Our main finding is that many recent empirical conclusions may be overly optimistic and fail to generalize under careful re-evaluation. We identify a surprising degree of sensitivity in LLM-based reasoning pipelines to seemingly minor design choices—ranging from decoding parameters, prompt formatting, and random seeds to the hardware and software stacks used during evaluation (see Table 1). Particularly concerning is the instability introduced by small benchmark sizes: for example, AIME’24 and AMC’23 each contain only 30–40 examples, making performance metrics highly volatile—where even one question can shift Pass@1 by over 3 percentage points. This leads to substantial variance across seeds, often resulting in double-digit performance swings that challenge the reliability of published results. In Section 3, we systematically analyze the root causes of this instability, including sampling variance, decoding configurations, evaluation frameworks, and hardware heterogeneity. We show that these factors can significantly distort conclusions if not carefully controlled.
@@
In Section 4, we propose a set of best practices aimed at improving reproducibility and rigor in reasoning benchmarks. We also re-evaluate recent techniques using a standardized and reproducible evaluation stack. Our findings are sobering—reinforcement learning (RL) applied to distillation-based models such as DeepSeek-R1 yields little to no statistically significant gains. Some methods, such as OpenRS, show promising results in original reports, but fail to hold up under repeated evaluation. RL training on base models like Qwen2.5 Math does show stronger performance, but still often underperforms instruction-tuned counterparts.111We note that OpenReasoner-Zero is a consistent exception, achieving competitive performance. Furthermore, RL-trained models exhibit significant performance drops on newer benchmarks such as AIME’25, echoing patterns of test set overfitting or “hill-climbing” observed in prior work (Golchin & Surdeanu, 2023; Roberts et al., 2023; Dominguez-Olmedo et al., 2024). In contrast, supervised fine-tuning (SFT) continues to deliver stable, generalizable improvements across benchmarks, underscoring its maturity as a training paradigm. These observations point to a critical need for more reliable and standardized evaluation protocols.
@@
2 Related Works
Language Model Reasoning (for Math). The recent releases of OpenAI-O1 (Jaech et al., 2024) (in September 2024), OpenAI-O3 (OpenAI, 2025) (in December 2024) and DeepSeek-R1 (DeepSeek-AI, 2025) (in January 2025), have spurred the language modelling community to work on improving the reasoning capabilites of language models. Several popular methods for improving those capabilites have emerged with supervised fine-tuning (SFT) and reinforcement learning (RL) being the two primary methods of interest (Uesato et al., 2022; Lightman et al., 2023; Lyu et al., 2025; Team, 2025). Recent works have built upon the DeepSeek-R1 recipe by proposing newer RL algorithms, including LCPO (Aggarwal & Welleck, 2025), REINFORCE++ (Hu, 2025), DAPO (Yu et al., 2025), DPO-VP (Tu et al., 2025), VinePPO (Kazemnejad et al., 2024), CPPO (Lin et al., 2025a), VAPO (Yue et al., 2025) and GRO (Cai, 2025). To gain a stronger understanding of how to induce mathematical capabilities, other works have conducted significant empirical studies exploring the design space of RL methods (Zeng et al., 2025b; Liu et al., 2025b; Team et al., 2025; Shao et al., 2024), including data scaling trends (Shen et al., 2025), curriculums (Wen et al., 2025b; Roux et al., 2025) and reward design (Gao et al., 2024a; Cui et al., 2025; Ma et al., 2023). Based on the success of these methods, there have also been recent efforts into scaling up reinforcement learning approaches to induce reasoning in domains beyond math, including code (Liu & Zhang, 2025; Xie et al., 2025; Jha et al., 2024; Yu et al., 2024), medicine (Zhang et al., 2025; Sim & Chen, 2024) and other sciences (Su et al., 2025; Yuan et al., 2025; Zeng et al., 2025a). Further, some works also explored scaling up RL-based approaches to modalities beyond just language, including vision (Ma et al., 2025; Meng et al., 2025; Huang et al., 2025; Peng et al., 2025; Chen et al., ; Deng et al., 2025; Liu et al., 2025c; Feng et al., 2025; Lin et al., 2025b). In our work, we objectively re-evaluate the claims made by several of these recent works under a standardized lens, and find that many of the reported gains do not hold up strongly when pitted on a level-playing field against well-tuned baselines.
Sobering Studies on ML Progress. Machine learning is a field of rapid progress. Due to the lightning speed of papers coming out across the various sub-fields of machine learning, practitioners and researchers often fail to rigorously evaluate algorithmic progress (Hutchinson et al., 2022; Dehghani et al., 2021; Machado et al., 2018; Ghosh et al., 2024; Balduzzi et al., 2018; Liao et al., 2021; Cawley & Talbot, 2010; Lipton & Steinhardt, 2019; Prabhu et al., 2024b). This has led to several papers showing that simple well-tuned baselines outperform months of progress on a specific sub-field in machine learning, including in continual learning (Prabhu et al., 2024a; 2020), active learning (Cawley, 2011) and test-time adaptation (Press et al., 2023). With the rapid influx of reasoning-based LMs, such statistically rigorous comparisons of models are ever more important—yet, despite the heavy use of RL-algorithms for driving progress in reasoning, there is very little mention of how different methods standardize their evaluations across different factors of variability. RL-algorithms themselves are known to be quite fickle to extremely minor variations including random seeds (Agarwal et al., 2021; Gorsane et al., 2022; Chan et al., 2019; Jordan et al., 2020; Patterson et al., 2024). Some works have even gone as far as suggesting that reliable benchmarking of RL-based methods is computationally infeasible (Jordan et al., 2024). Additionally, other works have demonstrated critical reliability issues in the generalization of frontier models to minor perturbations in the question inputs (Mirzadeh et al., 2024; Nezhurina et al., 2024; Srivastava et al., 2024), the type of tasks tested (Yan et al., 2025; Petrov et al., 2025; Dominguez-Olmedo et al., 2024; Roberts et al., 2025), metrics used (Liu et al., 2024) and in data-scarce scenarios (Udandarao et al., 2024; Kandpal et al., 2023; Parashar et al., 2024). Given such a volatile landscape, in this work, we aim to level the playing field across recent LM-methods that have been released and provide an objective look on the progress that the reasoning community has made. Our findings, which we discuss in the rest of the paper, are sobering at best.
3 Exploring the Design Space of Reasoning: What Matters Most?
Recent reasoning-focused language models are evaluated under highly heterogeneous conditions—including differences in evaluation frameworks and hardware, number of random seeds, temperature, and nucleus sampling parameters (top_p) (see Table 1). While prior work has examined the effect of sampling parameters in multiple-choice (Renze, 2024) and coding tasks (Arora et al., 2024), the influence of these choices remains underexplored for open-ended reasoning models—particularly those trained with reinforcement learning. In this section, we systematically assess how these evaluation design choices affect reported performance, and highlight the sources of variance that most impact the reliability of results.
3.1 Experimental Setup
We adopt a consistent experimental setup throughout this section, unless otherwise stated. Our analysis includes nine widely used models grouped into two commonly benchmarked size classes: 1.5B and 7B parameters. For the 1.5B class, we evaluate: DeepSeek-R1-Distill-1.5B (DeepSeek-AI, 2025), DeepScaleR-1.5B (Luo et al., 2025), II-1.5B-Preview (Intelligent Internet, 2025) , OpenRS1-1.5B, OpenRS2-1.5B, and OpenRS3-1.5B (Dang & Ngo, 2025). Note that DeepScaleR-1.5B, II-1.5B-Preview, and the OpenRS models are all initialized from DeepSeek-R1-Distill-1.5B and subsequently finetuned via reinforcement learning (e.g., GRPO (Shao et al., 2024)) to enhance mathematical reasoning capabilities. For the 7B class, we evaluate: DeepSeek-R1-Distill-7B, S1.1-7B (Muennighoff et al., 2025), and OpenThinker-7B (Team, 2025). Both S1.1-7B and OpenThinker-7B are finetuned Qwen2.5-7B-Instruct models (Yang et al., 2024a), trained using supervised learning on reasoning traces derived from DeepSeek-R1. All models are benchmarked on three widely used datasets: AIME’24 (AI-MO, ), AMC’23 (AI-MO, 2024), and MATH500 (Hendrycks et al., 2021), using the Pass@1 metric. Each result is averaged over multiple seeds and obtained on a standardized software stack (throguh a Docker image), and hardware with the following configuration: one 40 GB A100 GPU, an AMD 7302 32-core CPU, and 1TB RAM. All experiments were run using lighteval (Fourrier et al., 2023) with the vllm backend (Kwon et al., 2023).
@@ 3.2 Seed Variance in Evaluation
We begin by analyzing the variance induced purely by the random seed used during evaluation—an aspect often neglected in benchmarking practices. While recent work calls for statistical rigor (e.g., using error bars and multiple runs) (Bowyer et al., 2025; Biderman et al., 2024; Madaan et al., ), evaluations frequently rely on single-seed runs, obscuring potential variability. We assess the seed-induced variance across 20 independent evaluation runs for each of the nine models. Results are shown in Figure 2.
@@ 3.2.1 Can Bootstrapping Improve Mean Estimates?
To mitigate high variance, recent work has adopted bootstrapping—averaging multiple evaluation runs to stabilize results. For example, DeepSeek reports Pass@1 over 64 runs, while DeepScaleR uses 16. We study the effectiveness of this approach by bootstrapping estimates for AIME’24 using 1 to 10 evaluation runs. Figure 3 shows that while variance is extreme for and still large for , it reduces sharply for . Further analysis of variance across additional datasets is presented Figures 12 and 13.
@@ 3.2.2 Variance from Sampling Parameters: Temperature and top-p
Additionally, we investigate the impact of the temperature and top_p hyperparameter as prior works often employ different temperature and top_p settings when comparing the same model. To isolate the impact of varying temperature and top_p, we averaged pass@1 across seeds and compute variation of this estimate across temperature and top_p in a boxplot. Figure 6 and 7 show the performance variation. We see that temperature-induced and top_p-induced fluctuations not only affect performance estimates but also introduce substantial variability in performance itself, which can lead to unfair comparisons when evaluating the same model across different temperatures.
3.3 Variance from Hardware and Software Factors
Performance can also vary due to non-obvious factors like hardware and evaluation framework—yet this is rarely acknowledged. Models are often tested on heterogeneous systems and evaluated using different toolchains.For example, S-1.1 (Muennighoff et al., 2025) uses lm-evaluation-harness (Gao et al., 2024b), the OpenRS model suite uses lighteval (Fourrier et al., 2023), and II-1.5B-Preview uses evalscope (Alibaba ModelScope Community, ) for evaluation.

### v2

@@ Abstract
Reasoning has emerged as the next major frontier for language models (LMs), with rapid advances from both academic and industrial labs. However, this progress often outpaces methodological rigor, with many evaluations relying on benchmarking practices that lack transparency, robustness, or statistical grounding. In this work, we conduct a comprehensive empirical study and find that current mathematical reasoning benchmarks are highly sensitive to subtle implementation choices—including decoding parameters, random seeds, prompt formatting, and even hardware and software configurations. Performance gains reported in recent studies frequently hinge on unclear comparisons or unreported sources of variance. To address these issues, we propose a standardized evaluation framework with clearly defined best practices and reporting standards. Using this framework, we reassess recent methods and find that most reinforcement learning (RL) approaches yield only modest improvements—far below prior claims—and are prone to overfitting, especially on small-scale benchmarks like AIME’24. In contrast, supervised finetuning (SFT) methods show consistently stronger generalization in the settings we study. To foster reproducibility, we release all code, prompts, and model outputs, for reasoning benchmarks, establishing more rigorous foundations for future work.
@@ 1 Introduction
Reasoning has become central to recent advances in large language models (LLMs), playing a key role in nearly all frontier systems (Jaech et al., 2024; Anthropic, 2025; OpenAI, 2025a; xAI, 2025; Meta-AI, 2025; DeepMind, 2025). Recent months have seen a surge of research focused on understanding and improving LLM reasoning, accompanied by several open-source tools and training strategies (see Li et al. (2025c) for a survey). This momentum has sparked optimism that building capable, competitive reasoning models may soon be within reach.
@@
However, as evaluation practices shape the direction and perceived progress of the field (Liao et al., 2021), concerns around methodological rigor are growing. Non-reproducible or inconclusive evaluations can distort scientific understanding, misguide adoption, and skew future research priorities (Henderson et al., 2018; Marie et al., 2021; Musgrave et al., 2020; Prabhu et al., 2020; Andrychowicz et al., 2020; Colas et al., 2018). In the fast-moving area of LLM reasoning, where rapid publication cycles and benchmarking races are common, methodological shortcuts can quietly undermine progress. While concerns about reproducibility in LLM evaluations are well-documented (Reuel et al., 2024; Biderman et al., 2024), their persistence - especially in reasoning, calls for renewed scrutiny and higher standards.
@@
Motivated by a growing number of inconsistent empirical claims across the reasoning landscape, we conduct a rigorous investigation into the current state of reasoning benchmarks, focusing specifically on mathematical reasoning - one of the most widely used testbeds for evaluating algorithmic advances in this space (HuggingFaceH4, 2024; AI-MO, 2024a).
@@
Our main finding is that many recent empirical conclusions may be overly optimistic and fail to generalize under careful re-evaluation. We identify a surprising degree of sensitivity in LLM-based reasoning pipelines to seemingly minor design choices, ranging from decoding parameters, prompt formatting, and random seeds, to the hardware and software stacks used during evaluation (see Table 1). Particularly concerning is the instability introduced by small benchmark sizes: for example, AIME’24 and AMC’23 each contain only 30–40 examples, making performance metrics highly volatile, where even one question can shift Pass@1 by over 3 percentage points. This leads to substantial variance across seeds, often resulting in double-digit performance swings that challenge the reliability of published results. In Section 3, we systematically analyze the root causes of this instability, including sampling variance, decoding configurations, evaluation frameworks, and hardware heterogeneity. We show that these factors can significantly distort conclusions if not carefully controlled.
@@
In Section 4, we propose a set of best practices aimed at improving reproducibility and rigor in reasoning benchmarks. We also re-evaluate recent techniques using a standardized and reproducible evaluation stack. Our findings are sobering: reinforcement learning (RL) applied to distillation-based models such as DeepSeek-R1 yields little to no statistically significant gains. Some methods, such as OpenRS, show promising results in original reports, but fail to hold up under repeated evaluation. RL training on base models like Qwen2.5 Math does show stronger performance, but still often underperforms instruction-tuned counterparts.111We note that OpenReasoner-Zero is a consistent exception, achieving competitive performance. Furthermore, RL-trained models exhibit significant performance drops on newer benchmarks such as AIME’25, echoing patterns of test set overfitting or “hill-climbing” observed in prior work (Golchin & Surdeanu, 2023; Roberts et al., 2023; Dominguez-Olmedo et al., 2024). In contrast, supervised fine-tuning (SFT) continues to deliver stable, generalizable improvements across benchmarks, underscoring its maturity as a training paradigm. These observations point to a critical need for more reliable and standardized evaluation protocols.
@@
| L1-Exact | RL | DeepSeek-R1-Distill-1.5B | verl | 0.6 | 0.95 | 8,000 | – |
| L1-Max | RL | DeepSeek-R1-Distill-1.5B | verl | 0.6 | 0.95 | 8,000 | – |
@@
| SimpleRL-Zoo-7B | RL | Qwen2.5-7B | verl | 1 | 0.95 | 16,000 | 32 |
| SimpleRL-Zoo-Math-7B | RL | Qwen2.5-Math-7B | verl | 1 | 0.95 | 16,000 | 32 |
| OpenReasoner-Zero-7B | RL | Qwen2.7-7B | – | 1 | 1 | – | 16 |
| Eurus2 Prime | RL | Qwen2.5-Math-7B | custom | 0 | 1 | 4,096 | – |
2 Related Works
Language Model Reasoning (for Math). The recent releases of OpenAI-o1 (Jaech et al., 2024) (in September 2024), OpenAI-o3 (OpenAI, 2025a) (in December 2024) and DeepSeek-R1 (DeepSeek-AI, 2025) (in January 2025), have spurred the language modelling community to work on improving the reasoning capabilites of language models. Several popular methods for improving those capabilites have emerged with supervised fine-tuning (SFT) and reinforcement learning (RL) being the two primary methods of interest (Uesato et al., 2022; Lightman et al., 2023; Lyu et al., 2025; Open Thoughts, 2025). Recent works have built upon the DeepSeek-R1 recipe by proposing newer RL algorithms, including LCPO (Aggarwal & Welleck, 2025), REINFORCE++ (Hu, 2025), DAPO (Yu et al., 2025), DPO-VP (Tu et al., 2025), VinePPO (Kazemnejad et al., 2024), CPPO (Lin et al., 2025a), VAPO (Yue et al., 2025b) and GRO (Cai, 2025). To gain a stronger understanding of how to induce mathematical capabilities, other works have conducted significant empirical studies exploring the design space of RL methods (Zeng et al., 2025b; Liu et al., 2025c; Team et al., 2025; Shao et al., 2024), including data scaling trends (Shen et al., 2025), curriculums (Wen et al., 2025; Roux et al., 2025) and reward design (Gao et al., 2024a; Cui et al., 2025; Ma et al., 2023). Based on the success of these methods, there have also been recent efforts into scaling up reinforcement learning approaches to induce reasoning in domains beyond math, including code (Liu & Zhang, 2025; Xie et al., 2025; Jha et al., 2024; Yu et al., 2024), medicine (Zhang et al., 2025; Sim & Chen, 2024) and other sciences (Su et al., 2025; Yuan et al., 2025; Zeng et al., 2025a). Further, some works also explored scaling up RL-based approaches to modalities beyond just language, including vision (Ma et al., 2025; Meng et al., 2025; Huang et al., 2025; Peng et al., 2025; Chen et al., 2025; Deng et al., 2025; Liu et al., 2025d; Feng et al., 2025; Lin et al., 2025b). In our work, we objectively re-evaluate the claims made by several of these recent works under a standardized lens, and find that many of the reported gains do not hold up strongly when pitted on a level-playing field against well-tuned baselines.
Sobering Studies on ML Progress. Machine learning is a field of rapid progress. Due to the lightning speed of papers coming out across the various sub-fields of machine learning, practitioners and researchers often fail to rigorously evaluate algorithmic progress (Hutchinson et al., 2022; Dehghani et al., 2021; Machado et al., 2018; Ghosh et al., 2024; Balduzzi et al., 2018; Liao et al., 2021; Cawley & Talbot, 2010; Lipton & Steinhardt, 2019; Prabhu et al., 2024b; Card et al., 2020; Dror et al., 2018). This has led to several papers showing that simple well-tuned baselines outperform months of progress on a specific sub-field in machine learning, including in continual learning (Prabhu et al., 2024a; 2020), active learning (Cawley, 2011) and test-time adaptation (Press et al., 2023). With the rapid influx of reasoning-based LMs, such statistically rigorous comparisons of models are ever more important—yet, despite the heavy use of RL-algorithms for driving progress in reasoning, there is very little mention of how different methods standardize their evaluations across different factors of variability. RL-algorithms themselves are known to be quite fickle to extremely minor variations including random seeds (Agarwal et al., 2021; Gorsane et al., 2022; Chan et al., 2019; Jordan et al., 2020; Patterson et al., 2024). Some works have even gone as far as suggesting that reliable benchmarking of RL-based methods is computationally infeasible (Jordan et al., 2024). Additionally, other works have demonstrated critical reliability issues in the generalization of frontier models to minor perturbations in the question inputs (Mirzadeh et al., 2024; Nezhurina et al., 2024; Srivastava et al., 2024), the type of tasks tested (Yan et al., 2025b; Petrov et al., 2025; Dominguez-Olmedo et al., 2024; Roberts et al., 2025), metrics used (Liu et al., 2024) and in data-scarce scenarios (Udandarao et al., 2024; Kandpal et al., 2023; Parashar et al., 2024). Given such a volatile landscape, in this work, we aim to level the playing field across recent LM-methods that have been released and provide an objective look on the progress that the reasoning community has made. Our findings, which we discuss in the rest of the paper, are sobering at best.
3 Exploring the Design Space of Reasoning: What Matters Most?
Recent reasoning focused language models are evaluated under highly heterogeneous conditions—including differences in evaluation frameworks, hardware, number of random seeds, temperature, and nucleus sampling parameters (top_p) (see Table 1). While prior work has examined the effect of sampling parameters in multiple-choice (Renze, 2024) and coding tasks (Arora et al., 2024), the influence of these choices remains underexplored for open-ended reasoning models—particularly those trained with reinforcement learning. In this section, we systematically assess how these evaluation design choices affect reported performance, and highlight the sources of variance that most impact the reliability of results.
3.1 Experimental Setup
We adopt a consistent experimental setup throughout this section, unless otherwise stated. Our analysis includes nine widely used models grouped into two commonly benchmarked size classes: 1.5B and 7B parameters. For the 1.5B class, we evaluate: DeepSeek-R1-Distill-1.5B (DeepSeek-AI, 2025), DeepScaleR-1.5B (Luo et al., 2025), II-1.5B-Preview (Intelligent Internet, 2025) , OpenRS1-1.5B, OpenRS2-1.5B, and OpenRS3-1.5B (Dang & Ngo, 2025). Note that DeepScaleR-1.5B, II-1.5B-Preview, and the OpenRS models are all initialized from DeepSeek-R1-Distill-1.5B and subsequently finetuned via reinforcement learning (e.g., GRPO (Shao et al., 2024)) to enhance mathematical reasoning capabilities. For the 7B class, we evaluate: DeepSeek-R1-Distill-7B, S1.1-7B (Muennighoff et al., 2025), and OpenThinker-7B (Open Thoughts, 2025). Both S1.1-7B and OpenThinker-7B are finetuned Qwen2.5-7B-Instruct models (Qwen et al., 2025), trained using supervised learning on reasoning traces derived from DeepSeek-R1. All models are benchmarked on three widely used datasets: AIME’24 (AI-MO, 2024a), AMC’23 (AI-MO, 2024b), and MATH500 (Hendrycks et al., 2021), using the Pass@1 metric. Each result is averaged over multiple seeds and obtained on a standardized software stack (through a Docker image), and hardware with the following configuration: one 40 GB A100 GPU, an AMD 7302 32-core CPU, and 1TB RAM. All experiments were run using lighteval (Fourrier et al., 2023) with the vllm backend (Kwon et al., 2023).
@@ 3.2 Seed Variance in Evaluation
We begin by analyzing the variance induced purely by the random seed used during evaluation—an aspect often neglected in benchmarking practices. While recent work calls for statistical rigor (e.g., using error bars and multiple runs) (Bowyer et al., 2025; Biderman et al., 2024; Madaan et al., 2024), evaluations frequently rely on single-seed runs, obscuring potential variability. We assess the seed-induced variance across 20 independent evaluation runs for each of the nine models. Results are shown in Figure 2.
@@ 3.2.1 Can Bootstrapping Improve Mean Estimates?
To mitigate high variance, recent work has adopted bootstrapping: averaging multiple evaluation runs to stabilize results. For example, DeepSeek reports Pass@1 over 64 runs, while DeepScaleR uses 16. We study the effectiveness of this approach by bootstrapping estimates for AIME’24 using 1 to 60 evaluation runs. Figure 3 shows that while variance is extreme for and still large for , it reduces sharply for . Further analysis of variance across additional datasets is presented in Figures 14 and 16.
@@ 3.2.2 Variance from Sampling Parameters: Temperature and top-p
We additionally investigate the impact of the temperature and top_p hyperparameter as prior works often employ different temperature and top_p settings when comparing the same model. To isolate the impact of varying temperature and top_p, we averaged pass@1 across seeds and compute variation of this estimate across temperature and top_p in a boxplot. Figure 4 and 5 show the performance variation. We see that temperature-induced and top_p-induced fluctuations not only affect performance estimates but also introduce substantial variability in performance itself, which can lead to unfair comparisons when evaluating the same model across different temperatures.
@@
3.3 Effect of Prompt Format and Context Length
Prompt Format. Prompt formatting has a measurable impact on accuracy. As shown in Figure 9, models perform best when using math-specific prompts and their native chat templates. Omitting templates leads to performance drops, particularly for instruction-tuned models. We compare accuracy under three different prompt settings (see Table 6): (1) a math-specific prompt formatted using the model’s chat template, (2) only the model’s chat template with no additional prompt, and (3) no template at all, i.e., the question without any special tokens or instructions. Interestingly, while base models like Qwen2.5-Math may benefit from prompt-free setups (Liu et al., 2025c), instruction-tuned models rely heavily on format alignment. Thus, maintaining consistent and format-aware prompting is essential for maximizing instruction-tuned model performance.
3.4 Variance from Hardware & Software Factors: A Critical Consideration
Performance can also vary due to non-obvious factors like hardware and evaluation framework—yet this is rarely acknowledged. Models are often tested on heterogeneous systems and evaluated using different toolchains. For example, S-1.1 (Muennighoff et al., 2025) uses lm-evaluation-harness (Gao et al., 2024b), the OpenRS model suite uses lighteval (Fourrier et al., 2023), and II-1.5B-Preview uses evalscope (Alibaba ModelScope Community, 2025) for evaluation.